OpenStudy (anonymous):

Find the square root of 1,024 using either the factoring or extracting method.

OpenStudy (kropot72):

Start by dividing by small numbers that are factors. 2 is a good one. 1024/2 = 512 Divide by 2 again 512/2 = 256 and again 256/2 = 128 and again 128/2 = 64 and again 64/2 = 32 and again 32/2 = 16 and again 16/2 = 8 and again 8/2 = 4 and for the last time 4/2 = 2 Therefore\[1024=2^{10}\] \[\sqrt{1024}=\sqrt{2^{10}}=2^{\frac{10}{2}}=?\]

OpenStudy (kropot72):

\[2^{\frac{10}{2}}=2^{5}=32\]